# Magnetic field and kinetic helicity evolution in simulations of interacting disk galaxies

### v1

@@ 3 Numerical methods and overview of simulations
Figure 2: Time evolution of the separation $d$ between the centers of interacting galaxies for five cases. Sim. No.1 and 15 are central collisions ($\alpha_{\mathrm{b}}=0^{\circ}$) at different initial inclinations (see Table 2). For Sim. No. 3 ($\alpha_{\mathrm{b}}=20^{\circ}$) different interaction stages are visualized in Fig. 1. Sim. No. 6 and 18 are examples for grazing interactions at $\alpha_{\mathrm{b}}=45^{\circ}$. The grey dashed line marks a separation of 101010\;10 kpc, corresponding to the diameter of the center regions of the initial disks.
Figure 3: Evolution of galaxy shape parameters $s=b/a$ (left) and $q=c/a$ (right) of Sim. No. 1, 3, 6, which represent a sequence of simulations with increasing impact parameter for fixed initial inclinations.
@@ 4.1 Phenomenology of interacting disks
We begin with a phenomenological discussion of representative examples. In Fig. 1, density slices from an edge-on collision of two galaxies with $\alpha_{\mathrm{b}}=20^{\circ}$ (Sim. No. 3 in Table 2) are overlaid with a visualization of magnetic field lines using the line integral convolution method (LIC). [^4] Like iron filings surrounding a bar magnet, magnetic field lines are visualized by filtering a texture consisting of white-noise along local streamlines derived from the magnetic field data. The first three panels in Fig. 1 show the disks shortly before the collision, the time of closest approach (pericentric passage), and 100100100\,100Myr later. The separation of the galaxy centers is plotted as a function of time in Fig. 2. After reversing their motion, the galaxies are pulled toward each other once more and finally merge into a remnant of elliptical shape (Fig. 1, lower right). This is indicated by the time evolution of the shape parameters $s$ and $q$ plotted in Fig. 3. The two parameters specify the ratios of the three principal axes (thin disk: $s=1$, $q\sim 0$; sphere: $q=s=1$). As can be seen, the initially disk-like galaxies are strongly distorted during the interaction stages and evolve into an ellipsoid in the post-merger phase.
@@
The gas density indicated by the color map in Fig. 1 shows that a substantial amount of gas is ripped off the initial disks by tidal forces. The magnetic field structure reveals that the initially ordered toroidal fields inside the two disks evolves into highly turbulent structures expanding into the surroundings of the galaxies. After the first encounter, this can be mainly observed in the region in between the two disks (the so-called bridge) and in the tidal tails. The final remnant exhibits an unordered field that extends far into low-density regions. This suggests that interactions play a role in the magnetization of the CGM, although the magnetic field in a realistic cosmological environment is considerably larger than the weak background field of $10^{-25}\;$ assumed in our simulations.
@@
Figure 4: Slices of the gas density in combination with a line integral convolution of the magnetic field showing post-interaction stages for four different initial orientations. From left to right and top to bottom: Sim. No. 1, 3, 6, and 18 (see Table 2). While the disk inclinations are the same in the first three simulations, Sim. No. 6 and 18 share the impact parameter $\alpha_{\mathrm{b}}=45^{\circ}$.
@@
For nearly central collisions with a minimal separation below $10\;\mathrm{kpc}$, the disks are strongly perturbed after the first interaction (Fig. 4, upper panels). With increasing impact parameter, only the outer regions of the disks interact and less turbulence is produced. A bridge connecting the disks and tidal tails become the most prominent interaction features in the case of a grazing encounter (Fig. 4, lower left). Although the magnetic field remains mostly ordered at this stage, the disks eventually undergo substantial changes and finally merge, as indicated by the time evolution of the separation and shape parameters plotted in Figs. 1 and 3, respectively. The distortion of the disks is reduced further if angular momentum is aligned with velocity ($i_{1}=i_{2}=0^{\circ}$), i. e. the disk planes of both galaxies are oriented perpendicular to the velocity vector ${V}_{\mathrm{rel}}$ (Fig. 4, lower right). This type of collision is called face-on (group III in Table 2). Consequently, not only the impact parameter, but also the disk orientation has a significant influence on the outcome of galaxy interactions.
@@
Another example of a face-on collision from group III is shown in Fig. 5. In this simulation, the impact parameter is zero (Sim. No. 15). The galaxies have parallel angular momentum vectors, i. e., the rotation of the two disks is prograde. The morphology after the first encounter is markedly different from the central edge-on collision shown in Fig. 4 (upper left). Instead of tidal tails and a highly turbulent bridge, ring-like perturbations in the outer disks are accompanied by braid-like outward gas streams. As time progresses, more gas is scattered outwards and most of the outer disk disperses (Fig. 5, lower left). The remnant after about 1.51.51.5\;1.5Gyr is highly irregular and does not resemble a disk any longer (Fig. 5, lower right). In the following sections, we investigate whether the morphological differences between different types of interactions are reflected in magnetic field statistics.
@@
4.2 Evolution of average magnetic field
@@
Figure 6 shows the mean magnetic field strength in the primary galaxy as function of time for different subsets of simulations. The amplification of the magnetic field by interactions is indicated by the height of peaks. Edge-on collisions (group 0) with impact parameter varying between $\alpha_{\mathrm{b}}=0^{\circ}$ and $45^{\circ}$ are shown in the left plots. While the collision is central for $\alpha_{\mathrm{b}}=0^{\circ}$, the impact parameter $b$ is about 150150150\;150kpc for $\alpha_{\mathrm{b}}=45^{\circ}$. As expected from the phenomenological discussion in Sect. 4.1, the first maximum of the average magnetic field in center regions shows a decline with increasing impact parameter (Fig. 6, upper left). This indicates that the magnetic field in the inner disk is strongly amplified only for nearly central collisions. In contrast, differences between the average magnetic field in off-center regions are less pronounced because the outer disks are tidally disrupted even for relatively large impact parameters (Fig. 6, lower left). The largest field strength in an off-center region is observed for $\alpha_{\mathrm{b}}=15^{\circ}$ (Sim. No. 2). This suggests that the mixture of shock fronts and turbulence is particularly efficient in driving the amplification of the field if the collision is slightly off-center. During the second encounter, particularly strong fields are produced for $\alpha_{\mathrm{b}}$ in the range from $15^{\circ}$ and $\alpha_{\mathrm{b}}=25^{\circ}$ (Sim. No. 2 to 4). For $\alpha_{\mathrm{b}}\geq 25^{\circ}$, the maximum field strength in center regions is reached in the third peak. These trends show that in the case of gracing interactions several approaches are required to affect the magnetic field throughout the disks.
@@
Comparing simulations from different groups at a fixed impact parameter ($\alpha_{\mathrm{b}}=15^{\circ}$), we see that peaks approximately coincide, but the peak height varies with the disk orientation (right plots in Fig. 6). Peaks are particularly pronounced if the primary’s motion is parallel to its disk plane ($i_{1}=90^{\circ}$, groups 0, I, and II). For the face-on collision (Sim. No. 16 in group III), magnetic field amplification is significantly reduced both in center and off-center regions. The average magnetic field just prior to the first encounter is also smaller in this case. This is a consequence of stronger gas stripping if the disk moves face-on relative to the ambient medium, which removes magnetic energy from the cylindrical volume in which averages are computed. In groups IV and V, both disks are inclined with respect to their relative motion and their relative inclinations are $|i_{2}-i_{1}|=45^{\circ}$ and $90^{\circ}$, respectively. In these two cases, the peaks are comparable to the face-on collision, with the exception of the center region in Sim. No. 20 during the first encounter. Overall, our findings suggest that the most important parameter is the disk inclination, i. e., the angle between disk velocity and rotation axis, while the relative inclination $|i_{2}-i_{1}|$ also plays a role, but it is subdominant.
@@
Figure 7: Categorical Heatmaps displaying magnetic field strength at first interaction ordered by impact parameter angle $\alpha_{\mathrm{b}}$ and simulation group. Simulations are grouped by common initial inclinations of galaxies, i.e. analogous to Tab. 2 Group 0: 1-6, Group I: No. 7-10, Group II: 11-14, Group III: 15-18, Group IV: 19-22, Group V: 23-26. Shown are results for center (top) and off-center (bottom) regions at the first (left) and second (right) encounters. Parameter combinations that are not covered by our simulation suite or simulations that were not evolved over a sufficiently long time to reach the second interaction are left white.
@@
In virtually all cases, the peaks in the time evolution of the magnetic field strength are narrow (see Fig. 6), indicating that magnetic field amplification induced by interactions is transient. This behaviour can be explained by competing processes. As can be seen in Fig. 1, the magnetic field spreads over surrounding regions through tidal interactions, implying a gradual loss of magnetic field energy inside the disks. Field amplification by shock compression and turbulence acts against these losses. However, these processes are relatively short-lived.
@@ 4.3 Mean-field analysis
To analyse dynamo activity in our simulations, we decompose the numerically computed velocity and magnetic field into mean-field and turbulent components:
@@
We follow the approach of Ntormousi et al. (2020) and apply a mass-averaged filter with smoothing scale $\overline{\Delta}$ in every spatial dimension to compute the mean-field components. We choose $\overline{\Delta}=300\;$pc, corresponding to an average over $5^{3}$ grid cells (five-point stencil in one dimension). The resulting fluctuations, which are given by the difference between the local and mean fields, are associated with the smallest numerically resolved scales. While the smoothing scale applied by Ntormousi et al. (2020) is comparable to the scale height of a thin cooling disk and supernova-driven bubbles in the disk, our choice can be regarded as a compromise between constraining fluctuations to scales comparable to the grid resolution scale and the strong damping of numerical dissipation in this range of scales (see also Grete et al. 2017).
@@
Figure 8: Top row: Small-scale electromotive force $\mathcal{E}$ averaged over center regions for simulations of edge-on collisions (left) and simulations sharing the impact parameter $\alpha_{\mathrm{b}}=15^{\circ}$ (right). In each case, the EMF is normalized with respect to its value at $t=500\;$Myr. Bottom row: $\mathcal{E}$ relative to $\overline{\mathbf{v}}\times\overline{\mathbf{B}}$ (see eq. 17) averaged over center regions for the same simulations as in the top row. The corresponding magnetic field evolution is shown in Fig. 6.
Figure 8 shows the evolution of the small-scale EMF averaged over center regions for the same simulations as in Fig. 6. In each case, we normalize $\mathcal{E}$ by $\mathcal{E}_{500}$, i.e. the average vale at $t=500\;$Myr. At that point, the galaxies are still in the phase of the first approach, but perturbations in the initial disks have largely relaxed. Thereby, changes in $\mathcal{E}$ with respect to the state of the galaxies before the first encounter when $\mathcal{E}/\mathcal{E}_{500}\approx 1$ can be readily seen. Moreover, we concentrate the discussion on center regions. There is a steep rise of $\mathcal{E}$ in at the first encounter in nearly central collisions. In this phase, $\mathcal{E}/\mathcal{E}_{500}$ increases by up to two orders of magnitude. Only for larger impact parameters, the increase is smaller or even delayed until the second encounter (Sim. No. 5 and 6). In most simulations, the maximum small-scale EMF is reached around the second encounters. An exception is the central edge-on collisions (Sim. No. 1), where by far the strongest amplification occurs during the initial interaction. Apart from that, small-scale perturbations and, thus, turbulence tend to grow in subsequent interactions. However, this evolution is not sustained. In every simulation that was advanced for a sufficiently long period of time, we see that $\mathcal{E}$ gradually decreases in the final phase, when the galaxies have merged and small-scale flow is decaying. Although the averaged magnetic field is produced by both the mean-field and the small-scale EMF, its time evolution (Fig. 6, upper left and right) roughly reflects the evolution of $\mathcal{E}$. This suggests that $\mathcal{E}$ contributes significantly to the amplification of the magnetic field, particularly during close encounters. In the intermediate phases between encounters, $\mathcal{E}/\mathcal{E}_{500}\sim 10$ in most cases (see Fig. 8), while the average magnetic field decreases.
@@
which we evaluate within two adjacent cylindrical volumes that are immediately below ($K_{\mathrm{turb}}^{<}$) and above ($K_{\mathrm{turb}}^{>}$) the midplane, respectively (see also Ntormousi et al. 2020). With a thickness of 444\;4kpc and a radial extent of 202020\;20kpc, the combined volume above and below the midplane encompasses the center and offcenter regions for computing averages of the mangetic field strength. Since helicity is a pseudoscalar, reflection across the midplane reverses its sign (Yokoi 2013). Prior to interaction, $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ are comparable and have opposite signs. The time evolution is shown for selected simulations in Fig. 9. In the case of a central edge-on collision (Sim. No. 0), $K_{\mathrm{turb}}^{>}$ has a pronounced negative peak at the first encounter, which is roughly balanced by a positive peak of $K_{\mathrm{turb}}^{<}$. This phase is dominated by strong helical flows inside the tidal arms and the tidal bridge connecting the retracting collision partners. Assuming a linear relation between $\mathcal{E}$ and $K_{\mathrm{turb}}$ and an initially coherent toroidal mean field, the $\alpha$ effect preferentially amplifies the magnetic field in opposite directions above and below the midplane, which results in a quadrupolar field component (Ntormousi et al. 2020). After the first encounter, both $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ become small and irregular. For larger impact parameters (Fig. 9, Sim. No. 3 and 6), $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ remain antisymmetric in the post-interaction phase and throughout the second encounter. In the case of Sim. No. 3 ($\alpha_{\mathrm{b}}=20^{\circ}$), antisymmetry is eventually broken in the coalescence phase. A similar evolution is seen in Sim. No. 6 ($\alpha_{\mathrm{b}}=45^{\circ}$), but instead of the pronounced peaks at the second encounter, the average kinetic helicity decreases gradually at late time. Comparing to Fig. 8, pronounced antisymmetric peaks of $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ are associated with a decrease in $\mathcal{E}/(\overline{\@vec{v}}\times\overline{\@vec{B}})$. If the magnetic field is still more or less toroidal at that point, the resulting small-scale EMF tends to be antisymmetric and cancels out over the whole cylindrical region. The increasing EMF in post-interaction phases, on the other hand, indicates a growing influence of non-helical isotropic turbulence.
@@
Figure 9: Turbulent kinetic helicity defined by eq. (17) averaged over centered cylinders ($r\leq 5\;\mathrm{kpc}$, $|z|\leq 4\;\mathrm{kpc}$) below and above the galaxy disk midplane. Left Column: Simulations of edge-on collisions (group 0 with $\alpha_{\mathrm{b}}=0^{\circ}$, $20^{\circ}$, and $45^{\circ}$). Right column: Simulations sharing the impact parameter $\alpha_{\mathrm{b}}=15^{\circ}$ (groups II, III, and IV).
@@
For scenarios where the disk inclination differs from $90^{\circ}$, we see a markedly different evolution (Fig. 9, right column). As an example for face-on collisions (group III), Sim. No. 16 is shown. In this case, $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ have the same sign after the initial approach, i. e., the contributions from above and below the midplane are symmetric rather than antisymmetric. Since the two disks have parallel angular momentum, one would expect that the mean kinetic helicity of the lower half of one disk cancels the upper half of the other disk when they overlap. However, it appears that tidal interactions between the two disks fundamentally change their structure, resulting in a positive (first encounter) or negative (second encounter) net helicity. To a lesser degree this can also be seen in Sim. No. 20 (group IV, both disks are inclined), while Sim. No. 12 (group II, only secondary disk has inclination $0^{\circ}$) is a mixed case with alternating antisymmetric and symmetric episodes. These qualitative differences in the kinetic helicity evolution correspond well to differences in the small-scale EMF and the relatively weak magnetic field amplification in groups III to V, where both discs are inclined.

### v2

@@ 3 Numerical methods and overview of simulations
Figure 2: Time evolution of the separation, $d$, between the centers of interacting galaxies for five cases. Sim. No.1 and 15 are central collisions ($\alpha_{\mathrm{b}}=0^{\circ}$) at different initial inclinations (see Table 2). For Sim. No. 3 ($\alpha_{\mathrm{b}}=20^{\circ}$), different interaction stages are visualized in Fig. 1. Sim. No. 6 and 18 are examples for grazing interactions at $\alpha_{\mathrm{b}}=45^{\circ}$. The dashed gray line marks a separation of 101010\;10 kpc, corresponding to the diameter of the center regions of the initial disks.
Figure 3: Evolution of galaxy shape parameters $s=b/a$ (left) and $q=c/a$ (right) of Sim. No. 1, 3, and 6, which represent a sequence of simulations with increasing impact parameter for fixed initial inclinations.
@@ 4.1 Phenomenology of interacting disks
We begin with a phenomenological discussion of representative examples. In Fig. 1, density slices from an edge-on collision of two galaxies with $\alpha_{\mathrm{b}}=20^{\circ}$ (Sim. No. 3 in Table 2) are overlaid with a visualization of magnetic field lines using the line integral convolution method. [^5] Like iron filings surrounding a bar magnet, magnetic field lines are visualized by filtering a texture consisting of white-noise along local streamlines derived from the magnetic field data. The first three panels in Fig. 1 show the disks shortly before the collision, the time of closest approach (pericentric passage), and 100100100\,100Myr later. The separation of the galaxy centers is plotted as a function of time in Fig. 2. After reversing their motion, the galaxies are pulled toward each other once more and finally merge into a remnant of elliptical shape (Fig. 1, lower right). This is indicated by the time evolution of the shape parameters $s$ and $q$ plotted in Fig. 3. The two parameters specify the ratios of the three principal axes (thin disk: $s=1$, $q\sim 0$; sphere: $q=s=1$). As can be seen, the initially disk-like galaxies are strongly distorted during the interaction stages and evolve into an ellipsoid in the post-merger phase.
@@
The gas density indicated by the color map in Fig. 1 shows that a substantial amount of gas is ripped off the initial disks by tidal forces. The magnetic field structure reveals that the initially ordered toroidal fields inside the two disks evolves into highly turbulent structures expanding into the surroundings of the galaxies. After the first encounter, this can be mainly observed in the region in between the two disks (the so-called bridge) and in the tidal tails. The final remnant exhibits an unordered field that extends far into low-density regions. This suggests that interactions play a role in the magnetization of the circumgalactic medium, although the magnetic field in a realistic cosmological environment is considerably larger than the weak background field of $10^{-25}\;$ assumed in our simulations.
@@
Figure 4: Slices of the gas density in combination with a line integral convolution of the magnetic field showing post-interaction stages for four different initial orientations. From left to right and top to bottom: Sim. No. 1, 3, 6, and 18 (see Table 2). While the disk inclinations are the same in the first three simulations, Sim. No. 6 and 18 use the impact parameter $\alpha_{\mathrm{b}}=45^{\circ}$.
@@
For nearly central collisions with a minimal separation below $10\;\mathrm{kpc}$, the disks are strongly perturbed after the first interaction (Fig. 4, upper panels). With increasing impact parameter, only the outer regions of the disks interact and less turbulence is produced. A bridge connecting the disks and tidal tails become the most prominent interaction features in the case of a grazing encounter (Fig. 4, lower left). Although the magnetic field remains mostly ordered at this stage, the disks eventually undergo substantial changes and finally merge, as indicated by the time evolution of the separation and shape parameters plotted in Figs. 1 and 3, respectively. The distortion of the disks is reduced further if angular momentum is aligned with velocity ($i_{1}=i_{2}=0^{\circ}$), that is to say, if the disk planes of both galaxies are oriented perpendicular to the velocity vector ${V}_{\mathrm{rel}}$ (Fig. 4, lower right). This type of collision is called face-on (group III in Table 2). Consequently, not only the impact parameter, but also the disk orientation has a significant influence on the outcome of galaxy interactions.
@@
Another example of a face-on collision from group III is shown in Fig. 5. In this simulation, the impact parameter is zero (Sim. No. 15). The galaxies have parallel angular momentum vectors (i.e., the rotation of the two disks is prograde). The morphology after the first encounter is markedly different from the central edge-on collision shown in Fig. 4 (upper left). Instead of tidal tails and a highly turbulent bridge, ring-like perturbations in the outer disks are accompanied by braid-like outward gas streams. As time progresses, more gas is scattered outward and most of the outer disk disperses (Fig. 5, lower left). The remnant after about 1.51.51.5\;1.5Gyr is highly irregular and does not resemble a disk any longer (Fig. 5, lower right). In the following sections, we investigate whether the morphological differences between different types of interactions are reflected in magnetic field statistics.
@@
4.2 Average magnetic field evolution
@@
Figure 6 shows the mean magnetic field strength in the primary galaxy as a function of time for different subsets of simulations. The amplification of the magnetic field by interactions is indicated by the height of peaks. Edge-on collisions (group 0) with impact parameter varying between $\alpha_{\mathrm{b}}=0^{\circ}$ and $45^{\circ}$ are shown in the left plots. While the collision is central for $\alpha_{\mathrm{b}}=0^{\circ}$, the impact parameter $b$ is about 150150150\;150kpc for $\alpha_{\mathrm{b}}=45^{\circ}$. As expected from the phenomenological discussion in Sect. 4.1, the first maximum of the average magnetic field in center regions shows a decline with increasing impact parameter (Fig. 6, upper left). This indicates that the magnetic field in the inner disk is strongly amplified only for nearly central collisions. In contrast, differences between the average magnetic field in off-center regions are less pronounced because the outer disks are tidally disrupted even for relatively large impact parameters (Fig. 6, lower left). The largest field strength in an off-center region is observed for $\alpha_{\mathrm{b}}=15^{\circ}$ (Sim. No. 2). This suggests that the mixture of shock fronts and turbulence is particularly efficient in driving the amplification of the field if the collision is slightly off-center. During the second encounter, particularly strong fields are produced for $\alpha_{\mathrm{b}}$ in the range from $15^{\circ}$ and $\alpha_{\mathrm{b}}=25^{\circ}$ (Sim. No. 2 to 4). For $\alpha_{\mathrm{b}}\geq 25^{\circ}$, the maximum field strength in center regions is reached in the third peak. These trends show that in the case of gracing interactions several approaches are required to affect the magnetic field throughout the disks.
@@
Comparing simulations from different groups at a fixed impact parameter ($\alpha_{\mathrm{b}}=15^{\circ}$), we see that peaks approximately coincide, but the peak height varies with the disk orientation (right plots in Fig. 6). Peaks are particularly pronounced if the primary’s motion is parallel to its disk plane ($i_{1}=90^{\circ}$, groups 0, I, and II). For the face-on collision (Sim. No. 16 in group III), magnetic field amplification is significantly reduced both in center and off-center regions. The average magnetic field just prior to the first encounter is also smaller in this case. This is a consequence of stronger gas stripping if the disk moves face-on relative to the ambient medium, which removes magnetic energy from the cylindrical volume in which averages are computed. In groups IV and V, both disks are inclined with respect to their relative motion and their relative inclinations are $|i_{2}-i_{1}|=45^{\circ}$ and $90^{\circ}$, respectively. In these two cases, the peaks are comparable to the face-on collision, with the exception of the center region in Sim. No. 20 during the first encounter. Overall, our findings suggest that the most important parameter is the disk inclination (i.e., the angle between the disk velocity and rotation axis); the relative inclination $|i_{2}-i_{1}|$ also plays a role, but it is subdominant.
@@
Figure 7: Categorical heat maps displaying the magnetic field strength at the first interaction, ordered by impact parameter angle ($\alpha_{\mathrm{b}}$) and simulation group. Simulations are grouped by common initial inclinations of galaxies, i.e., analogous to Table 2 Group 0: 1-6, Group I: No. 7-10, Group II: 11-14, Group III: 15-18, Group IV: 19-22, Group V: 23-26. Shown are results for center (top) and off-center (bottom) regions at the first (left) and second (right) encounters. Parameter combinations that are not covered by our simulation suite or simulations that were not evolved over a sufficiently long time to reach the second interaction are left white.
@@
In virtually all cases, the peaks in the time evolution of the magnetic field strength are narrow (see Fig. 6), indicating that magnetic field amplification induced by interactions is transient. This behavior can be explained by competing processes. As can be seen in Fig. 1, the magnetic field spreads over surrounding regions through tidal interactions, implying a gradual loss of magnetic field energy inside the disks. Field amplification by shock compression and turbulence acts against these losses. However, these processes are relatively short-lived.
@@ 4.3 Mean-field analysis
To analyze dynamo activity in our simulations, we decomposed the numerically computed velocity and magnetic field into mean-field and turbulent components:
@@
We followed the approach of Ntormousi et al. (2020) and applied a mass-averaged filter with smoothing scale $\overline{\Delta}$ in every spatial dimension to compute the mean-field components. We chose $\overline{\Delta}=300\;$pc, which corresponds to an average over $5^{3}$ grid cells (five-point stencil in one dimension). The resulting fluctuations, which are given by the difference between the local and mean fields, are associated with the smallest numerically resolved scales. While the smoothing scale applied by Ntormousi et al. (2020) is comparable to the scale height of a thin cooling disk and supernova-driven bubbles in the disk, our choice can be regarded as a compromise between constraining fluctuations to scales comparable to the grid resolution scale and the strong damping of numerical dissipation in this range of scales (see also Grete et al. 2017).
@@
Figure 8: Small-scale EMF, $\mathcal{E,}$ averaged over center regions for simulations of edge-on collisions (left) and simulations with the impact parameter $\alpha_{\mathrm{b}}=15^{\circ}$ (right). Top row: EMF normalized with respect to its value at $t=500\;$Myr. Bottom row: $\mathcal{E}$ relative to $\overline{\mathbf{v}}\times\overline{\mathbf{B}}$ (see Eq. 17) averaged over center regions for the same simulations as in the top row. The corresponding magnetic field evolution is shown in Fig. 6.
Figure 8 shows the evolution of the small-scale EMF averaged over center regions for the same simulations as in Fig. 6. In each case, we normalized $\mathcal{E}$ by $\mathcal{E}_{500}$ (i.e., the average vale at $t=500\;$Myr). At that point, the galaxies are still in the phase of the first approach, but perturbations in the initial disks have largely relaxed. Thereby, changes in $\mathcal{E}$ with respect to the state of the galaxies before the first encounter when $\mathcal{E}/\mathcal{E}_{500}\approx 1$ can be readily seen. Moreover, we concentrate the discussion on center regions. There is a steep rise of $\mathcal{E}$ in at the first encounter in nearly central collisions. In this phase, $\mathcal{E}/\mathcal{E}_{500}$ increases by up to two orders of magnitude. Only for larger impact parameters, the increase is smaller or even delayed until the second encounter (Sim. No. 5 and 6). In most simulations, the maximum small-scale EMF is reached around the second encounters. An exception is the central edge-on collisions (Sim. No. 1), where by far the strongest amplification occurs during the initial interaction. Apart from that, small-scale perturbations and, thus, turbulence tend to grow in subsequent interactions. However, this evolution is not sustained. In every simulation that was advanced for a sufficiently long period of time, we see that $\mathcal{E}$ gradually decreases in the final phase, when the galaxies have merged and small-scale flow is decaying. Although the averaged magnetic field is produced by both the mean-field and the small-scale EMF, its time evolution (Fig. 6, upper left and right) roughly reflects the evolution of $\mathcal{E}$. This suggests that $\mathcal{E}$ contributes significantly to the amplification of the magnetic field, particularly during close encounters. In the intermediate phases between encounters, $\mathcal{E}/\mathcal{E}_{500}\sim 10$ in most cases (see Fig. 8), while the average magnetic field decreases.
@@
which we evaluated within two adjacent cylindrical volumes that are immediately below ($K_{\mathrm{turb}}^{<}$) and above ($K_{\mathrm{turb}}^{>}$) the midplane, respectively (see also Ntormousi et al. 2020). With a thickness of 444\;4kpc and a radial extent of 202020\;20kpc, the combined volume above and below the midplane encompasses the center and off-center regions for computing averages of the magnetic field strength. Since helicity is a pseudoscalar, reflection across the midplane reverses its sign (Yokoi 2013). Prior to interaction, $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ are comparable and have opposite signs. The time evolution is shown for selected simulations in Fig. 9. In the case of a central edge-on collision (Sim. No. 0), $K_{\mathrm{turb}}^{>}$ has a pronounced negative peak at the first encounter, which is roughly balanced by a positive peak of $K_{\mathrm{turb}}^{<}$. This phase is dominated by strong helical flows inside the tidal arms and the tidal bridge connecting the retracting collision partners. Assuming a linear relation between $\mathcal{E}$ and $K_{\mathrm{turb}}$ and an initially coherent toroidal mean field, the $\alpha$ effect preferentially amplifies the magnetic field in opposite directions above and below the midplane, which results in a quadrupolar field component (Ntormousi et al. 2020). After the first encounter, both $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ become small and irregular. For larger impact parameters (Fig. 9, Sim. No. 3 and 6), $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ remain antisymmetric in the post-interaction phase and throughout the second encounter. In the case of Sim. No. 3 ($\alpha_{\mathrm{b}}=20^{\circ}$), antisymmetry is eventually broken in the coalescence phase. A similar evolution is seen in Sim. No. 6 ($\alpha_{\mathrm{b}}=45^{\circ}$), but instead of the pronounced peaks at the second encounter, the average kinetic helicity decreases gradually at late time. Comparing to Fig. 8, pronounced antisymmetric peaks of $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ are associated with a decrease in $\mathcal{E}/(\overline{\@vec{v}}\times\overline{\@vec{B}})$. If the magnetic field is still more or less toroidal at that point, the resulting small-scale EMF tends to be antisymmetric and cancels out over the whole cylindrical region. The increasing EMF in post-interaction phases, on the other hand, indicates a growing influence of non-helical isotropic turbulence.
@@
Figure 9: Turbulent kinetic helicity defined by Eq. (17) averaged over centered cylinders ($r\leq 5\;\mathrm{kpc}$, $|z|\leq 4\;\mathrm{kpc}$) below and above the galaxy disk midplane. Left column: Simulations of edge-on collisions (group 0 with $\alpha_{\mathrm{b}}=0^{\circ}$, $20^{\circ}$, and $45^{\circ}$). Right column: Simulations with the impact parameter $\alpha_{\mathrm{b}}=15^{\circ}$ (groups II, III, and IV).
@@
For scenarios where the disk inclination differs from $90^{\circ}$, we see a markedly different evolution (Fig. 9, right column). As an example for face-on collisions (group III), Sim. No. 16 is shown. In this case, $K_{\mathrm{turb}}^{>}$ and $K_{\mathrm{turb}}^{<}$ have the same sign after the initial approach, meaning that the contributions from above and below the midplane are symmetric rather than antisymmetric. Since the two disks have parallel angular momentum, one would expect that the mean kinetic helicity of the lower half of one disk cancels the upper half of the other disk when they overlap. However, it appears that tidal interactions between the two disks fundamentally change their structure, resulting in a positive (first encounter) or negative (second encounter) net helicity. To a lesser degree this can also be seen in Sim. No. 20 (group IV, both disks are inclined), while Sim. No. 12 (group II, only secondary disk has inclination $0^{\circ}$) is a mixed case with alternating antisymmetric and symmetric episodes. These qualitative differences in the kinetic helicity evolution correspond well to differences in the small-scale EMF and the relatively weak magnetic field amplification in groups III to V, where both disks are inclined.